OpenStudy (anonymous):

x-4 x + 4 ------ = ------- x+4 x + 1 solve the equation for x

OpenStudy (anonymous):

Cross Multiply

OpenStudy (anonymous):

(x-4)(x+1)=(x+4)^2 can you solve it from here?

OpenStudy (anonymous):

x^2+x-4x-4=x^2+8x+16 -3x-4=8x+16 11x=-20 x=...?

OpenStudy (anonymous):

x=-20/11

OpenStudy (anonymous):

sorry i was answering history questions i got that answer as well ty